# MODE: Efficient Time Series Prediction with Mamba Enhanced by Low-Rank Neural ODEs

### v1

@@ Abstract
Time series prediction plays a pivotal role across diverse domains such as finance, healthcare, energy systems, and environmental modeling. However, existing approaches often struggle to balance efficiency, scalability, and accuracy, particularly when handling long-range dependencies and irregularly sampled data. To address these challenges, we propose MODE, a unified framework that integrates Low-Rank Neural Ordinary Differential Equations (Neural ODEs) with an Enhanced Mamba architecture. As illustrated in our framework, the input sequence is first transformed by a Linear Tokenization Layer and then processed through multiple Mamba Encoder blocks, each equipped with an Enhanced Mamba Layer that employs Causal Convolution, SiLU activation, and a Low-Rank Neural ODE enhancement to efficiently capture temporal dynamics. This low-rank formulation reduces computational overhead while maintaining expressive power. Furthermore, a segmented selective scanning mechanism, inspired by pseudo-ODE dynamics, adaptively focuses on salient subsequences to improve scalability and long-range sequence modeling. Extensive experiments on benchmark datasets demonstrate that MODE surpasses existing baselines in both predictive accuracy and computational efficiency. Overall, our contributions include: (1) a unified and efficient architecture for long-term time series modeling, (2) integration of Mamba’s selective scanning with low-rank Neural ODEs for enhanced temporal representation, and (3) substantial improvements in efficiency and scalability enabled by low-rank approximation and dynamic selective scanning. The code and data used in this work are available at the following link https://github.com/XsChen524/mode-mamba-ts.
@@ I Introduction
Time series prediction Lim and Zohren (2021); Wen et al. (2022); Zhang et al. (2025c); Esling and Agon (2012); ZHANG et al. (2018) is a fundamental task in machine learning and statistics, underpinning a wide range of real-world applications, including finance, healthcare, climate modeling, social networks Zheng et al. (2023, 2024a, 2024b); Zhang et al. (2025f); Zheng et al. (2025), and urban forecasting Zhang et al. (2025a, 2024, e); ZHANG et al. ; Mahalakshmi et al. (2016). The overarching objective is to learn temporal dependencies within sequential data in order to accurately forecast future observations. Despite significant progress, time series prediction remains challenging due to the complex characteristics of real-world data, which often exhibit nonlinear dynamics, long-range dependencies, and irregular sampling patterns.
Traditional approaches, such as autoregressive models Chen et al. (2011) and recurrent neural networks (RNNs) Medsker et al. (2001), typically struggle to capture long-term dependencies or complex temporal behaviors. Recent advances, particularly those based on Transformers Vaswani et al. (2017) and state-space models (SSMs) Auger-Méthé et al. (2021), have shown enhanced modeling capability for sequential data. However, these models are often limited by high computational costs, insufficient scalability, and difficulty in handling irregularly sampled time series effectively.
@@ II-A1 Mamba
The Mamba framework (Gu and Dao, 2023) defines a sequence-to-sequence transformation using four main parameters $(\mathbf{A},\mathbf{B},\mathbf{C},\Delta)$ in a continuous-time state-space form:
@@
which inherits useful properties from the underlying continuous system, such as resolution invariance (Nguyen et al., 2022). A central innovation of Mamba is that the parameters $\mathbf{B}$, $\mathbf{C}$, and the step size $\Delta$ are made input-dependent, so that the model can dynamically adapt its dynamics to the characteristics of the sequence.
@@ III-A Experiment Settings
Datasets. To rigorously assess the effectiveness of our proposed model, we curated a diverse collection of 9 real-world datasets from established sources (Zhou et al., 2023, 2021) for comprehensive evaluation. These datasets span multiple domains, including electricity consumption, and the 4 ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2), among others. Widely recognized in the research community, these datasets are instrumental in addressing challenges across various fields, such as transportation analysis and energy management. Detailed statistical information for each dataset is provided in Table I.
@@ III-B Detailed Baseline Descriptions
We evaluate MODE against tencontemporary approaches representing three major architectural paradigms: Transformer-based (6 methods), MLP-based (3 methods), and state-space model (SSM) based (1 method) techniques. The Transformer-based baselines include Autoformer (Wu et al., 2021), which fuses time series decomposition with an autocorrelation module to capture periodic temporal dependencies while avoiding conventional attention operations; FEDformer (Zhou et al., 2022), which embeds spectral-domain transformations into the attention mechanism to improve computational efficiency while maintaining a global receptive field; Crossformer (Zhang and Yan, 2022), which applies multidimensional attention over patched temporal segments, where patching enhances local feature extraction but its effectiveness diminishes for extremely long input sequences; DLinear (Zeng et al., 2023), which shows that two simple linear layers acting on decomposed trend and residual components can outperform many sophisticated attention-based models across a range of forecasting tasks; PatchTST (Huang et al., 2024), which adopts segmented time-step embeddings with channel-wise isolated processing to support the discovery of temporal patterns across multiple scales; and iTransformer (Liu et al., 2023), which reverses the conventional attention setup to focus on cross-variate interactions, though its MLP-based sequence-flattening tokenization is less effective at modeling hierarchical temporal structures. The MLP-based approaches comprise TimesNet (Wu et al., 2022), which converts univariate time series into two-dimensional periodic maps to enable unified modeling of both within-cycle and across-cycle patterns; RLinear (Li et al., 2023), which introduces reversible normalization in a channel-independent linear design and sets new baselines for models that rely solely on linear operations; and TiDE (Das et al., 2023), which leverages stacked fully connected layers in an encoder–decoder layout to capture temporal structure while entirely forgoing recurrence and attention. Finally, the state-space model variant S-Mamba (Wang et al., 2025) uses per-variate tokenization together with bidirectional Mamba layers to model cross-variate relationships, further supported by feedforward networks for temporal dynamics.
@@ III-C Effectiveness
Table II reports the experimental results of MODE and baselines on eight common time series forecasting benchmarks: ETTm1, ETTm2, ETTh1, ETTh2, Electricity, Exchange, Weather, Solar-Energy, and PEMS08. The lookback window $L$ is fixed at 96, and the forecasting horizon $T$ is 96, 192, 336, or 720. We report Mean Squared Error (MSE) and Mean Absolute Error (MAE), with the best and second-best results in bold and underline, respectively.
Across almost all datasets and horizons, MODE yields the best or second-best performance. On average, it attains the lowest MSE and MAE, surpassing strong baselines such as S-Mamba, iTransformer, PatchTST, and FEDformer. In particular, MODE reaches an average MSE of 0.178 and MAE of 0.265, corresponding to improvements of about 11.2% in MSE and 9.0% in MAE over the strongest baseline.
These gains hold across diverse settings: on ETT (industrial load) MODE improves long-horizon predictions; on large-scale datasets (Electricity, Exchange) it better captures long-term dependencies; and on challenging datasets (Weather, Solar-Energy) it preserves low errors and stable performance across all horizons. The effectiveness and robustness of MODE stem from three components: (1) continuous-time low-rank Neural ODEs for irregular sampling, (2) a selective scanning mechanism that focuses computation on informative segments, and (3) integration with the Mamba architecture for efficient, scalable temporal representations, jointly enabling state-of-the-art accuracy with high efficiency.
TABLE II: We present comprehensive results of MODE and baselines on the ETTh1, ETTh2, Electricity, Exchange, Weather, and Solar-Energy datasets. The lookback length $L$ is fixed at 96, and the forecast length $T$ varies across 96, 192, 336, and 720. Bold font denotes the best model and underline denotes the second best. All baseline results are obtained from Wang et al. (2025).
@@ III-D Ablation Study
To assess the contribution of each component in MODE, we conduct ablation studies summarized in Table III. We analyze: (1) the use of Neural ODEs, (2) dynamic vs. static low-rank parameterization, (3) the effect of rank size $r$, and (4) the interaction between the ODE formulation and the Mamba architecture. Experiments are run on ETTm1, Weather, and ECL with forecast horizons $T\in\{96,192,336,720\}$, using MSE and MAE.
Effect of Low-Rank Neural ODEs. Compared with standard Mamba and S-Mamba, our Neural ODE variants consistently improve performance across datasets and horizons. For example, at horizon 96 on ETTm1, MODE with static low-rank ODE achieves an MSE of 0.324, outperforming Standard Mamba (0.326) and S-Mamba (0.320). This shows that continuous-time dynamics help capture irregular temporal patterns that discrete models miss.
Dynamic versus Static Low-Rank Parameterization. The dynamic variant further improves over the static one. On Weather at horizon 96, Dynamic Low-Rank ODE achieves the best MSE/MAE (0.167/0.210), surpassing the static ODE (0.170/0.213). This suggests that adaptive ranks better capture time-varying dependencies, improving stability and accuracy.
@@
Influence of Architectural Components. We also test variants that replace Mamba enhanced by low rank ODE with vanilla Mamba, S-Mamba block, attention + FFN, or pure linear layers. Attention solution shows advantages in short-term forecasting, for its ability to capture high correlation between future and recent values, and respond to local variations by assigning higher weights to nearby timesteps. But MODE shows competitive performance with a considerable number of best or 2nd best error results in all prediction length, indicating that both the ODE formulation and its integration into the Mamba pipeline are key to MODE’s forecasting gains.
@@ III-E Efficiency Comparison
Fig. 3 illustrates the efficiency comparison between our proposed MODE and several strong baselines, including S-Mamba, PatchTST, iTransformer, Flowformer, Informer, BiMamba4T, Transformer, and DLinear, on the Weather and ETTm1 datasets. The x-axis represents the average training time per iteration, and the y-axis shows the Mean Squared Error (MSE), while the bubble size indicates GPU memory consumption. As shown, MODE achieves an excellent balance between accuracy and efficiency, yielding lower MSE with competitive training time and small memory footprint. Specifically, MODE achieves comparable or better accuracy than PatchTST and iTransformer while requiring significantly less training time and reduced memory usage (only 0.076 GB on ETTm1 and 0.126 GB on Weather). In contrast, Transformer-based and Flowformer architectures incur higher computational costs and memory overhead. These results demonstrate that the low-rank Neural ODE formulation and selective scanning mechanism enable MODE to achieve both high predictive accuracy and superior computational efficiency.
@@ III-F Robustness Study
To evaluate the robustness of MODE against input perturbations, we conduct experiments on the ETTm1 dataset by injecting Gaussian noise with varying standard deviations into the input sequences. Fig. 4 presents the results under four forecasting horizons $T\in\{96,192,336,720\}$. The left axis reports the absolute error values (MSE and MAE), while the right axis shows the relative performance degradation as noise intensity increases. As illustrated, MODE maintains stable performance under low to moderate noise levels (standard deviation $\leq 0.3$), with only marginal increases in MSE and MAE. Even under heavy noise injection ($\text{std}=0.5$), the error growth remains moderate compared to typical Transformer-based models. This robustness arises from the continuous-time dynamics of the Neural ODE formulation, which smooths noisy variations, and the selective scanning mechanism, which emphasizes informative temporal segments. These results confirm that MODE is resilient and reliable in noisy and uncertain environments.
@@ III-H Case Study
Accurate Trend Capture: Across all scenarios, MODE (blue solid line) consistently follows the ground truth patterns more closely than baseline methods. This is particularly evident in the U-shaped temperature curves observed throughout different days, where MODE accurately captures both the magnitude and timing of temperature variations.
@@
Figure 6: Case Study.
@@ IV Related Work
Modern time series forecasting has progressed rapidly with the development of deep sequence models, particularly Transformer-based architectures Vaswani et al. (2017). These models capture long-range temporal dependencies through the self-attention mechanism, enabling accurate multistep forecasting. Extensions of the Transformer framework Lim and Zohren (2021); Torres et al. (2021); Zheng et al. (2021) incorporate causal masking to preserve the temporal order, while hierarchical designs such as Pyraformer Liu et al. (2021) and cross-variate modules like Crossformer Zhang and Yan (2022) further enhance contextual representation. However, their quadratic time complexity $\mathcal{O}(N^{2})$ and memory cost limit scalability to long sequences. To address this, numerous studies have explored linearized or window-based attention mechanisms, yet these often trade off precision for efficiency. Large-scale pretraining frameworks such as Moirai Woo et al. (2024) offer another direction but differ substantially in setup and scale from our controlled evaluation.
Recently, Mamba-based architectures Gu and Dao (2023); Zhang et al. (2025d, f, b, c) have emerged as promising state-space models, demonstrating strong performance in temporal reasoning and long-context sequence modeling. While these methods improve computational efficiency, they remain limited in handling irregularly sampled sequences and continuous-time dynamics.

### v2

@@ Abstract
Time series prediction plays a pivotal role across diverse domains such as finance, healthcare, energy systems, and environmental modeling. However, existing approaches often struggle to balance efficiency, scalability, and accuracy, particularly when handling long-range dependencies and irregularly sampled data. To address these challenges, we propose MODE, a unified framework that integrates Low-Rank Neural Ordinary Differential Equations (Neural ODEs) with an Enhanced Mamba architecture. As illustrated in our framework, the input sequence is first transformed by a Linear Tokenization Layer and then processed through multiple Mamba Encoder blocks, each equipped with an Enhanced Mamba Layer that employs Causal Convolution, SiLU activation, and a Low-Rank Neural ODE enhancement to efficiently capture temporal dynamics. This low-rank formulation reduces computational overhead while maintaining expressive power. Furthermore, a segmented selective scanning mechanism, inspired by pseudo-ODE dynamics, adaptively focuses on salient subsequences to improve scalability and long-range sequence modeling. Extensive experiments on benchmark datasets demonstrate that MODE surpasses existing baselines in both predictive accuracy and computational efficiency. Overall, our contributions include: (1) a unified and efficient architecture for long-term time series modeling, (2) integration of Mamba’s selective scanning with low-rank Neural ODEs for enhanced temporal representation, and (3) substantial improvements in efficiency and scalability enabled by low-rank approximation and dynamic selective scanning. The code and data used in this work are available at the following link: https://github.com/XsChen524/mode-mamba-ts.
@@ I Introduction
Time series prediction lim2021time ; wen2022transformers ; zhang2025autohformer ; esling2012time ; zhang2018time is a fundamental task in machine learning and statistics, underpinning a wide range of real-world applications, including finance, healthcare, climate modeling, social networks zheng2023cipl ; zheng2024hycorec ; zheng2024mitigating ; zhang2025fldmamba ; zheng2025cirec , and urban forecasting zhang2025efficient ; zhang2024survey ; zhang2025survey ; zhangprompts ; mahalakshmi2016survey . The overarching objective is to learn temporal dependencies within sequential data in order to accurately forecast future observations. Despite significant progress, time series prediction remains challenging due to the complex characteristics of real-world data, which often exhibit nonlinear dynamics, long-range dependencies, and irregular sampling patterns.
Traditional approaches, such as autoregressive models ARIMA2 and recurrent neural networks (RNNs) medsker2001recurrent , typically struggle to capture long-term dependencies or complex temporal behaviors. Recent advances, particularly those based on Transformers vaswani2017attention and state-space models (SSMs) auger2021guide , have shown enhanced modeling capability for sequential data. However, these models are often limited by high computational costs, insufficient scalability, and difficulty in handling irregularly sampled time series effectively.
@@ II-A1 Mamba
The Mamba framework (gu2023mamba, ) defines a sequence-to-sequence transformation using four main parameters $(\mathbf{A},\mathbf{B},\mathbf{C},\Delta)$ in a continuous-time state-space form:
@@
which inherits useful properties from the underlying continuous system, such as resolution invariance (nguyen2022s4nd, ). A central innovation of Mamba is that the parameters $\mathbf{B}$, $\mathbf{C}$, and the step size $\Delta$ are made input-dependent, so that the model can dynamically adapt its dynamics to the characteristics of the sequence.
@@ III-A Experiment Settings
Datasets. To rigorously assess the effectiveness of our proposed model, we curated a diverse collection of 9 real-world datasets from established sources (haoyietal-informerEx-2023, ; haoyietal-informer-2021, ) for comprehensive evaluation. These datasets span multiple domains, including electricity consumption, and the 4 ETT datasets (ETTh1, ETTh2, ETTm1, ETTm2), among others. Widely recognized in the research community, these datasets are instrumental in addressing challenges across various fields, such as transportation analysis and energy management. Detailed statistical information for each dataset is provided in Table I.
@@ III-B Detailed Baseline Descriptions
We evaluate MODE against tencontemporary approaches representing three major architectural paradigms: Transformer-based (6 methods), MLP-based (3 methods), and state-space model (SSM) based (1 method) techniques. The Transformer-based baselines include Autoformer (wu2021autoformer, ), which fuses time series decomposition with an autocorrelation module to capture periodic temporal dependencies while avoiding conventional attention operations; FEDformer (zhou2022fedformer, ), which embeds spectral-domain transformations into the attention mechanism to improve computational efficiency while maintaining a global receptive field; Crossformer (zhang2022crossformer, ), which applies multidimensional attention over patched temporal segments, where patching enhances local feature extraction but its effectiveness diminishes for extremely long input sequences; DLinear (zeng2023transformers, ), which shows that two simple linear layers acting on decomposed trend and residual components can outperform many sophisticated attention-based models across a range of forecasting tasks; PatchTST (huang2024long, ), which adopts segmented time-step embeddings with channel-wise isolated processing to support the discovery of temporal patterns across multiple scales; and iTransformer (liu2023itransformer, ), which reverses the conventional attention setup to focus on cross-variate interactions, though its MLP-based sequence-flattening tokenization is less effective at modeling hierarchical temporal structures. The MLP-based approaches comprise TimesNet (wu2022timesnet, ), which converts univariate time series into two-dimensional periodic maps to enable unified modeling of both within-cycle and across-cycle patterns; RLinear (li2023revisiting, ), which introduces reversible normalization in a channel-independent linear design and sets new baselines for models that rely solely on linear operations; and TiDE (das2023long, ), which leverages stacked fully connected layers in an encoder–decoder layout to capture temporal structure while entirely forgoing recurrence and attention. Finally, the state-space model variant S-Mamba (wang2025mamba, ) uses per-variate tokenization together with bidirectional Mamba layers to model cross-variate relationships, further supported by feedforward networks for temporal dynamics.
@@ III-C Effectiveness
Table II reports the experimental results of MODE and baselines on eight common time series forecasting benchmarks: ETTm1, ETTm2, ETTh1, ETTh2, Electricity, Exchange, Weather, Solar-Energy, and PEMS08. The lookback window $L$ is fixed at 96, and the forecasting horizon $T$ is 96, 192, 336, or 720. We report Mean Squared Error (MSE) and Mean Absolute Error (MAE) as evaluating metrics respectively.
Given baseline from wang2024mamba and experimental results of MODE under optimal hyperparameter configurations, it yields the best or second-best performance across almost all datasets and horizons. On average, it attains the lowest MSE and MAE, surpassing strong baselines such as S-Mamba, iTransformer, PatchTST, and FEDformer.
These gains hold across diverse settings: on ETT (industrial load) MODE improves short-term predictions; on large-scale datasets (Electricity, Exchange) it better captures long-term dependencies; and on challenging datasets (Weather, Solar-Energy) it preserves low errors and stable performance across all horizons. The effectiveness and robustness of MODE stem from three components: (1) continuous-time low-rank Neural ODEs for irregular sampling, (2) a selective scanning mechanism that focuses computation on informative segments, and (3) integration with the Mamba architecture for efficient, scalable temporal representations, jointly enabling state-of-the-art accuracy with high efficiency.
TABLE II: We present comprehensive results of MODE and baselines on the ETTh1, ETTh2, Electricity, Exchange, Weather, and Solar-Energy datasets. The lookback length $L$ is fixed at 96, and the forecast length $T$ varies across 96, 192, 336, and 720. Bold font denotes the best model and underline denotes the second best. All baseline results are obtained from wang2025mamba .
@@ III-D Ablation Study
To assess the contribution of each component in MODE, we conduct ablation studies summarized in Table III. We analyze: (1) the use of Neural ODEs, (2) dynamic vs. static low-rank parameterization, (3) the effect of rank size $r$, and (4) the interaction between the ODE formulation and the Mamba architecture. Experiments are run on ETTm1, Weather, and ECL with forecast horizons $T\in\{96,192,336,720\}$, using evaluating metrics MSE and MAE.
Effect of Low-Rank Neural ODEs. Compared with vanilla Mamba and bidirectional Mamba blocks in parallel, Neural ODE variants are on par with or outperform models without ODE implementations across datasets and horizons. For example, at horizon 96 on ETTm1, MODE with static low-rank ODE achieves an MSE of 0.324, outperforming vanilla Mamba (0.326) and bidirectional Mamba (0.331) and is close to attention mechanism. It also achieves best or second best performances in horizon 336 and 720 experiments. This shows that continuous-time dynamics help capture irregular temporal patterns that discrete models miss.
Dynamic versus Static Low-Rank Parameterization. The dynamic variant further improves over the static one. On Weather at horizon 96, Dynamic Low-Rank ODE achieves the best MSE/MAE (0.167/0.210), surpassing the static ODE (0.170/0.213). This suggests that, at the expense of some time and memory efficiency, adaptive ranks better capture time-varying dependencies, improving stability and accuracy.
@@
Influence of Architectural Components. We also test variants that replace Mamba enhanced by low rank ODE with vanilla Mamba, S-Mamba block, attention + FFN, or pure linear layers. Attention solution shows advantages in short-term forecasting, for its ability to capture high correlation between future and recent values, and respond to local variations by assigning higher weights to nearby timesteps. But MODE shows competitive performance with a considerable number of best or second best error results in all prediction length, indicating that both the ODE formulation and its integration into the Mamba pipeline are key to MODE’s improvements.
@@ III-E Efficiency Comparison
Fig. 3 illustrates the efficiency comparison between our proposed MODE and several strong baselines, including S-Mamba, PatchTST, iTransformer, Flowformer, Informer, Transformer, and DLinear, on the Weather and ETTm1 datasets. The x-axis represents the average training time per iteration, the y-axis shows MSE results, and the bubble size indicates allocated GPU memory under same batching settings (128 for PatchTST, 16 for other models). MODE with static low-rank ODE achieves an excellent balance between accuracy and efficiency, yielding lower MSE with competitive training time and small memory footprint. Specifically, MODE achieves comparable or better accuracy than PatchTST and iTransformer while requiring significantly less training time and reduced memory usage (only 0.037 GB on ETTm1 and 0.137 GB on Weather). In contrast, Transformer-based and Flowformer architectures incur higher computational costs and memory overhead. These results demonstrate that the low-rank Neural ODE formulation and selective scanning mechanism enable MODE to achieve both high predictive accuracy and superior computational efficiency.
@@ III-F Robustness Study
To evaluate the robustness of MODE against input perturbations, we conduct experiments on the ETTm1 dataset by injecting Gaussian noise with varying standard deviations into the input sequences. Fig. 4 presents the results under four forecasting horizons $T\in\{96,192,336,720\}$. The left axis reports the absolute error values (MSE and MAE), while the right axis shows the relative performance degradation as noise intensity increases. As illustrated, MODE maintains stable performance under low to moderate noise levels (standard deviation $\leq 0.3$), with only marginal increases in MSE and MAE. Even under heavy noise injection ($\text{std}=0.5$), the error growth is accelerated but remains relatively moderate. Except for horizon 96 experiments, MSE and MAE growths are consistently kept within 10%. This robustness arises from the continuous-time dynamics of the Neural ODE formulation, which smooths noisy variations, and the selective scanning mechanism, which emphasizes informative temporal segments. These results confirm that MODE is resilient and reliable in noisy and uncertain environments.
@@ III-H Case Study
Accurate Trend Capture: Across all scenarios, MODE (purple solid line) consistently follows the ground truth patterns more closely than baseline methods. This is particularly evident in the U-shaped temperature curves observed throughout different days, where MODE accurately captures both the magnitude and timing of temperature variations.
@@
Figure 6: Case study of MODE on ETTm1 across different horizons and scenarios.
@@ IV Related Work
Modern time series forecasting has progressed rapidly with the development of deep sequence models, particularly Transformer-based architectures vaswani2017attention . These models capture long-range temporal dependencies through the self-attention mechanism, enabling accurate multistep forecasting. Extensions of the Transformer framework lim2021time ; torres2021deep ; zheng2021graph incorporate causal masking to preserve the temporal order, while hierarchical designs such as Pyraformer liu2021pyraformer and cross-variate modules like Crossformer zhang2022crossformer further enhance contextual representation. However, their quadratic time complexity $\mathcal{O}(N^{2})$ and memory cost limit scalability to long sequences. To address this, numerous studies have explored linearized or window-based attention mechanisms, yet these often trade off precision for efficiency. Large-scale pretraining frameworks such as Moirai woo2024unified offer another direction but differ substantially in setup and scale from our controlled evaluation.
Recently, Mamba-based architectures gu2023mamba ; zhang2025hmamba ; zhang2025fldmamba ; zhang2025m2rec ; zhang2025autohformer have emerged as promising state-space models, demonstrating strong performance in temporal reasoning and long-context sequence modeling. While these methods improve computational efficiency, they remain limited in handling irregularly sampled sequences and continuous-time dynamics.